Boyd Company has a line of credit with State Bank. Boyd can borrow up to $400,000 at any time over the course of the 2018 calendar year. The following table shows the prime rate expressed as an annual percentage along with the amounts borrowed and repaid during 2018. Boyd agreed to pay interest at an annual rate equal to 2 percent above the bank’s prime rate. Funds are borrowed or repaid on the first day of each month. Interest is payable in cash on the last day of the month. The interest rate is applied to the outstanding monthly balance. For example, Boyd pays 7 percent (5 percent + 2 percent) annual interest on $100,000 for the month of January.

Month Amount Borrowed or (Repaid) Prime Rate for the Month
January $ 100,000 5 %
February 70,000 6
March (30,000 ) 7
April through October No change No change
November (50,000 ) 6
December (40,000 ) 5

Boyd earned $45,000 of cash revenue during 2018.

Required

  1. Show the effects of these transactions on the financial statements using a horizontal statements model like the one shown here. Use a + to indicate increase, − for decrease, and if the element is not affected, leave the cell blank. In the Cash Flow column, indicate whether the item is an operating activity (OA), investing activity (IA), or financing activity (FA).
  2. What is the total amount of interest expense paid for 2018?
BOYD CO.
Horizontal Statements Model
Date Balance Sheet Income Statement Statement of Cash Flow
Assets = Liabilities + Equity Revenue Expense = Net Income
1/1 = + =
1/31 = + =
2/1 = + =
2/28 = + =
3/1 = + =
3/31 = + =
4/30 = + =
5/31 = + =
6/30 = + =
7/31 = + =
8/31 = + =
9/30 = + =
10/31 = + =
11/1 = + =
11/30 = + =
12/1 = + =
12/31 = + =

NOTE: For answers Assets and Statement of cash flow second tab are answered with FA, IA, or OA. The rest of the tabs are answered with +, -, or +/-
Answer #1
Date Balance Sheet income Statement Statement of Cash Flow
Assets = Liabilities + Equity Revenue - Expenses = Net Income
              45,000 = +                    45,000                    45,000 - =                45,000 OA
01-Jan           1,00,000 =                 1,00,000 +                            -                              -   -                     -   =                         -   FA
Jan-31           1,44,417 =                 1,00,000 +                    44,417                            -   -                  583 =                    -583 FA
01-Feb               70,000 =                    70,000 +                            -                              -   - =                         -   FA
28-Feb           2,13,283 =                 1,70,000 +                    43,283                            -   -               1,133 =                 -1,133 FA
01-Mar             -30,000 =                  -30,000 +                            -                              -   - =                         -   FA
Mar-31           1,82,233 =                 1,40,000 +                    42,233                            -   -               1,050 =                 -1,050 FA
01-Apr                       -   =                             -   +                            -                              -   - =                         -   FA
30-Apr           1,81,183 =                 1,40,000 +                    41,183                            -   -               1,050 =                 -1,050 FA
01-May                       -   =                             -   +                            -                              -   - =                         -   FA
31-May           1,80,133 =                 1,40,000 +                    40,133                            -   -               1,050 =                 -1,050 FA
01-Jun                       -   =                             -   +                            -                              -   - =                         -   FA
30-Jun           1,79,083 =                 1,40,000 +                    39,083                            -   -               1,050 =                 -1,050 FA
01-Jul                       -   =                             -   +                            -                              -   - =                         -   FA
31-Jul           1,78,033 =                 1,40,000 +                    38,033                            -   -               1,050 =                 -1,050 FA
01-Aug                       -   =                             -   +                            -                              -   - =                         -   FA
31-Aug           1,76,983 =                 1,40,000 +                    36,983                            -   -               1,050 =                 -1,050 FA
01-Sep                       -   =                             -   +                            -                              -   - =                         -   FA
30-Sep           1,75,933 =                 1,40,000 +                    35,933                            -   -               1,050 =                 -1,050 FA
01-Oct                       -   =                             -   +                            -                              -   - =                         -   FA
31-Oct           1,74,883 =                 1,40,000 +                    34,883                            -   -               1,050 =                 -1,050 FA
01-Nov             -50,000 =                  -50,000 +                            -                              -   - =                         -   FA
30-Nov           1,24,283 =                    90,000 +                    34,283                            -   -                  600 =                    -600 FA
01-Dec             -40,000 =                  -40,000 +                            -                              -   - =                         -   FA
31-Dec               83,992 =                    50,000 +                    33,992                            -   -                  292 =                    -292 FA
Ending Balance 83992 50000 33992 45000            11,008                33,992

Since date of revenue earned is not mentioned it has been taken on first day

Interest calculation
Jan-31                 583 =100000*(5+2)%/12
28-Feb             1,133 =170000*(6+2)%/12
Mar-31             1,050 =140000*(7+2)%/12
30-Apr             1,050 =140000*(7+2)%/12
31-May             1,050 =140000*(7+2)%/12
30-Jun             1,050 =140000*(7+2)%/12
31-Jul             1,050 =140000*(7+2)%/12
31-Aug             1,050 =140000*(7+2)%/12
30-Sep             1,050 =140000*(7+2)%/12
31-Oct             1,050 =140000*(7+2)%/12
30-Nov                 600 =90000*(6+2)%/12
31-Dec                 292 =50000*(5+2)%/12

2. Total Interest expense paid for 2018 = $11008.33